# Short-wavelength mesophases in the ground states of core-softened particles in two-dimensions

### v2

@@ 3.1 Phase diagram
The systematic determination of the ground-state phase diagram was carried out over a wide range of densities $\rho\in[0.7,2.5]$ and values of the hard-core parameter $C\in[10^{-4},10^{-0.5}]$, expressed more naturally as $\ell_{C}\equiv-\log C$. Figure 3 summarizes the phase behavior of the system under these parameters. Several important observations emerge from this diagram.
@@
To begin with, the overall structure of the phase diagram confirms that the addition of a hard-core repulsion to the GEM-4 ultra-soft interaction strongly enriches the phase behavior, suppressing the otherwise continuous sequence of cluster occupancies characteristic of the pure GEM-4 models [5, 45, 47]. Indeed, looking at the smallest values of $C$ explored (highest values of $\ell_{C}$), i.e., for weakest hard-core repulsion, one observes a sequence of cluster phases with occupancy numbers that grows with increasing density. This is expected from the known behavior of cluster crystal phases in systems with pure ultra-soft potentials, but with two remarkable differences. The first being the presence of more than one phase associated with a given occupation number. The second is that, as it will be discussed in detail shortly, the oblique lattice formed by the center of the clusters in some cases is closer to a triangular lattice than in others, depending on the occupancy number.
@@ 3.2 Ground-state configurations
As discussed previously, cluster anisotropy distorts the lattice, even in the regime when the size of the clusters is relatively small in comparison with the lattice spacing of the cluster lattice. This behavior is however much stronger for more elongated clusters, such as when we have dimer phases. In Fig. 5, we explore the behavior of the equilibrium $\theta$ angle for different values of $\ell_{C}$ and $\rho$, characterizing how close to the triangular case ($\theta=60^{\circ}$) the general oblique cluster lattice is for the different cluster phases. As can be observed, for the Cluster 3 or 3a phases, where the internal symmetry of the clusters are compatible with the triangular lattice, the lattice angles are very close to $60^{\circ}$. On the other hand, for the Cluster 2 or 2a phases, the distortion is larger and converges very slowly to $60^{\circ}$ as $C\rightarrow 0$ ($\ell_{C}\to\infty$). This slow convergence indicate, as mentioned before, that even for quite small clusters the equilibrium cluster lattice is significantly anisotropic.
@@ 3.3 Coexistence regions
where $\varepsilon\equiv E/N$ stands for the energy per particle function, which is determined from the ground state energy calculation. Coexistence limits are then obtained by solving the following coupled equations
@@
In Figure 6 we show the resulting phase diagram once we take into account the presence of coexistence regions. As can be observed this improvement modify significantly the thermodynamic phase diagram of the system, suppressing or eliminating entirely many of the pure phases in our system. This is the case, for instance, of the Cluster 2, Cluster 2a and Cluster 3a phases which presents significant coexistence regions at their phase boundaries. Moreover, we can verify the thermodynamic consistency of our calculations by examining the continuity of the extension of the coexistence regions when bifurcations in these regions occurs varying $\ell_{C}$. It is worth noticing that the extension of a coexistence region is an indicator of how strong a first order transition is, or in other words, how different the curves $\varepsilon(\rho)$ behave for each phase in the vicinity of the crossing density. Interestingly, the lower the values of $C$ the stronger are the first order transitions developed by the model. This indicates that the discontinuous transitions between configurations with different cluster occupancies are stronger than those in which single particle lattices reorganize in a different structure.
To complement the results for the phase diagram presented we show in Fig. 7 the behavior of the energy per particle, pressure and chemical potential increasing density at specific values of $\ell_{C}$. A closer look at this figure reveals the characteristic plateaus in the pressure and chemical potential expected in first order transitions. These plateaus are produced due to the presence of phase-separated domains, where the system locally minimizes its energy while maintaining mechanical and chemical equilibrium. Notably, the widths of most coexistence regions ($\Delta\rho$) are significant, suggesting that the transitions, while discontinuous, do not require extremely fine density control for experimental realization.
@@
Physically, the coexistence region tends to appear between phases with strongly mismatched local motifs, where continuous parameter deformations are energetically costly or forbidden. For example, coexistence is prominent between Cluster 2 and Cluster 3a, or between Cluster 3 and Cluster 4, whereas transitions involving Oblique and Rectangular structures have narrower coexistence intervals due to the possibility of smooth lattice deformations.
@@ 4 Conclusions
In this work, we have explored the ground-state phase diagram of a two-dimensional core-softened particle system, defined by the combination of the GEM-4 ultra-soft potential and an inverse power-law interaction with a decay exponent of 6 acting as a hard-core repulsion. At first, we conducted an exploration of the phase diagram using both molecular dynamics simulations and numerical energy minimization based on a general ansatz for $n$-occupied cluster phases with an independent orientational order in up to two sublattices. This strategy enabled us to identify the short-wavelength mesophases typically present in this model and to formulate a set of specific ansatz for each phase, respecting their symmetries while relying on the minimal set of parameters and constraints needed for their proper description.
@@
Beyond the cluster phases, as the hard-core repulsion becomes stronger, we have also identified single-particle phases, including triangular and square lattice arrangements, as well as phases associated with stripe-like mesophases (rectangular and oblique lattices) and hole-like mesophases (kagome and honeycomb lattices). The construction of the specific ansatz and the minimization of their energies not only allowed us to characterize and localize these phases in the phase diagram, but also to delineate coexistence regions near the phase transitions, thus mapping and classifying them as either discontinuous or continuous.
The detailed description of the short-wavelength region of the ground-state phase diagram presented in this work can establish the basis for future investigations of thermal and quantum fluctuation effects on mesophases. For example, this system could be an excellent candidate for studying the melting of anisotropic two-dimensional crystalline phases [71, 72], or, in the context of interacting bosons, for characterizing the corresponding quantum ground states [7, 32, 33, 73, 35]. Moreover, the classical behavior described here could be extended to higher densities, where longer-wavelength mesophases [3] may give rise to even more complex phenomena than those reported in the present study.

### v3

@@ 3.1 Phase diagram
The systematic determination of the ground-state phase diagram was carried out over a wide range of densities $\rho\in[0.7,2.5]$ and values of the hard-core strength parameter $\ell_{C}\in[0.5,4.0]$. Figure 3 summarizes the phase behavior of the system under these parameters. Several important observations emerge from this diagram.
@@
To begin with, the overall structure of the phase diagram confirms that the addition of a hard-core repulsion to the GEM-4 ultra-soft interaction strongly enriches the phase behavior, suppressing the otherwise contiguous sequence of cluster occupancies characteristic of the pure GEM-4 models [5, 45, 47]. Indeed, looking at the smallest values of $C$ explored (highest values of $\ell_{C}$), i.e., for weakest hard-core repulsion, one observes a sequence of cluster phases with occupancy numbers that grows with increasing density. This is expected from the known behavior of cluster crystal phases in systems with pure ultra-soft potentials, but with two remarkable differences. The first being the presence of more than one phase associated with a given occupation number. The second is that, as it will be discussed in detail shortly, the oblique lattice formed by the center of the clusters in some cases is closer to a triangular lattice than in others, depending on the occupancy number.
@@ 3.2 Ground-state configurations
As discussed previously, cluster anisotropy distorts the lattice, even in the regime when the size of the clusters is relatively small in comparison with the lattice spacing of the cluster lattice. This behavior is however much stronger for more elongated clusters, such as when we have dimer phases. In Fig. 6, we explore the behavior of the equilibrium $\theta$ angle for different values of $\ell_{C}$ and $\rho$, characterizing how close to the triangular case ($\theta=60^{\circ}$) the general oblique cluster lattice is for the different cluster phases. As can be observed, for the Cluster 3 or 3a phases, where the internal symmetry of the clusters are compatible with the triangular lattice, the lattice angles are very close to $60^{\circ}$. On the other hand, for the Cluster 2 or 2a phases, the distortion is larger and converges very slowly to $60^{\circ}$ as $C\rightarrow 0$ ($\ell_{C}\to\infty$). This slow convergence indicate, as mentioned before, that even for quite small clusters the equilibrium cluster lattice is significantly anisotropic. It is worth mentioning that this effect is noticeable in the structure factors of the simulation configurations shown in Fig. 1, indicating that the trimer phases retains its six-fold rotational symmetry more effectively, whereas in dimer phases this symmetry is typically reduced to two-fold.
@@ 3.3 Coexistence regions
where $\varepsilon\equiv E/N$ stands for the energy per particle function, which is determined from the ground state energy calculation. To obtain the pressure and chemical potential, which require the derivative of $\varepsilon(\rho)$, we use a cubic spline interpolation of the discrete energy data. The coexistence boundaries $(\rho_{1},\rho_{2})$ are then found by numerically solving the following coupled equations using a two-dimensional bisection method [72]:
@@
In Figure 7 we show the resulting phase diagram once we take into account the presence of coexistence regions. As can be observed this improvement modifies significantly the thermodynamic phase diagram of the system, suppressing or eliminating entirely many of the pure phases in our system. This is the case, for instance, of the Cluster 2, Cluster 2a and Cluster 3a phases which presents significant coexistence regions at their phase boundaries. Moreover, we can verify the thermodynamic consistency of our calculations by examining the continuity of the extension of the coexistence regions when bifurcations in these regions occurs varying $\ell_{C}$. It is worth noticing that the extension of a coexistence region is an indicator of how strong a first order transition is, or in other words, how different the curves $\varepsilon(\rho)$ behave for each phase in the vicinity of the crossing density. Interestingly, the lower the values of $C$ the stronger are the first order transitions developed by the model. This indicates that the discontinuous transitions between configurations with different cluster occupancies are stronger than those in which single particle lattices reorganize in a different structure.
To complement the results for the phase diagram presented we show in Fig. 8 the behavior of the energy per particle, pressure and chemical potential increasing density at specific values of $\ell_{C}$. The energy per particle as a function of density $\epsilon(\rho)$ is given by the lower envelope of the curves shown on the left column of this figure, where the colored curves correspond to pure phases and the gray curve corresponds to the energy per particle of the coexistence of neighboring phases given by eq. (12). Values of $\epsilon$ higher than the lower envelope indicate metastability of pure phases. A closer look at this figure reveals the characteristic plateaus in the pressure and chemical potential expected in first order transitions. These plateaus are produced due to the presence of phase-separated domains, where the system locally minimizes its energy while maintaining mechanical and chemical equilibrium. Notably, the widths of most coexistence regions ($\Delta\rho$) are significant, suggesting that the transitions, while discontinuous, do not require extremely fine density control for experimental realization.
@@
Physically, the coexistence region tends to appear between phases with strongly mismatched local motifs, where continuous parameter deformations are energetically costly or forbidden. For example, coexistence is prominent between Cluster 2 and Cluster 3a, or between Cluster 3 and Cluster 4, whereas transitions involving Oblique and Rectangular structures have narrower coexistence intervals due to the possibility of smooth lattice deformations. Thermodynamically, continuous transitions correspond to points where the energy-per-particle curves of the two phases meet with the same first derivative with respect to density, as can be seen for instance in Fig. 8 in the transition between Cluster-2a and Rectangular phases in the case $\ell_{C}=2.51$. Consequently, the Maxwell construction applied in such transition would consistently yield a coexistence region of zero width.
3.4 Quasicrystalline ground-states
In order to present a more complete picture of the phase diagram presented in Fig. 4, in this section we address the cases where our variational analysis of the ground-states could not capture the behavior observed in our simulations. As reported in previous studies [25, 26, 24], two-dimensional systems with HCSS potentials frequently presents quasicrystalline phases in their phase diagrams. In our case, we observe that the simulations discussed in section 2 produce low temperature configurations with decagonal and dodecagonal rotational symmetries, as shown in Fig. 9. Such configurations are observed in regions of the phase diagram that exhibit a high degree of frustration. The decagonal quasicrystalline configurations are observed in a region near the phase boundaries of the kagome phase with both triangular and honeycomb phases, whereas the dodecagonal quasicrystalline configurations are observed in a lower density region, near the phase boundaries between honeycomb, rectangular and square lattice phases. These regions are highlighted in the ground-state phase diagram shown in Fig. 10.
[…]
Figure 9: Decagonal (a) and dodecagonal (b) low temperature molecular dynamics configurations for systems with $N=900$ particles, together with the corresponding values of the density, relative strength of the hard-core interaction and temperature. The associated structure factors of the configurations are shown in the insets confirming their 10-fold or 12-fold rotational symmetry.
Figure 10: Phase diagram corrected by molecular dynamics simulations. The highlighted areas are a guide to the eye for the regions where our simulation grid identified the decagonal (10-QC) and dodecagonal (12-QC) quasicrystalline phases. Dots indicate the actual positions where simulations converged to quasicrystalline configurations.
To further validate our characterization of the low-temperature phases, we compared the potential energies obtained from Langevin simulations with those resulting from the minimization of the variational ansatz. For all periodic phases captured by the ansatz, the agreement is excellent, with energy differences well below the numerical resolution of the simulations. Interestingly, the quasicrystalline phases observed in the frustrated region exhibit energies that are extremely close to those of the corresponding minimized ansatz — typically differing by $\mathscr{O}(10^{-3})$. The quasicrystals display slightly lower energies, suggesting that aperiodic order may outperform the periodic candidates proposed in the variational analysis. This reinforces the view that competing length scales can stabilize nearly degenerate structures.
@@ 4 Conclusions
In this work, we have explored the ground-state phase diagram of a two-dimensional core-softened particle system, defined by the combination of the GEM-4 ultra-soft potential and an inverse power-law interaction with a decay exponent of 6 acting as a hard-core repulsion. At first, we conducted an exploration of the phase diagram using both molecular dynamics simulations and numerical energy minimization based on a general ansatz for $n$-occupied cluster phases with an independent orientational order in up to two sublattices. This strategy enabled us to identify the short-wavelength mesophases typically present in this model and to formulate a set of specific ansatz for each phase, respecting their symmetries while relying on the minimal set of parameters and constraints needed for their proper description. While our ansatz set was guided by two exploratory techniques and previously reported phases in similar models, we are aware that more complex periodic phases were observed in similar systems [26, 24]. Nevertheless, we are confident that, in this sense, our results captures the relevant physics of this system.
@@
Beyond the cluster phases, as the hard-core repulsion becomes stronger, we have also identified single-particle phases, including triangular and square lattice arrangements, as well as phases associated with stripe-like mesophases (rectangular and oblique lattices) and hole-like mesophases (kagome and honeycomb lattices). The construction of the specific ansatz and the minimization of their energies not only allowed us to characterize and localize these phases in the phase diagram, but also to calculate the coexistence regions corresponding to the first order phase transitions. Moreover, the fact that we are always locating local energy minima, rather than merely stationary points in the periodic configuration space, already provides strong evidence of the mechanical stability of the obtained phases. Furthermore, since all reported phases were also observed in simulations following careful equilibration and annealing, we can confirm their mechanical stability from a fully general perspective.
The rich phase diagram is confirmed through molecular dynamics simulation using a combination of full inertial Langevin dynamics, parallel tempering, and low-temperature annealing. These simulations were essential to prove the existence of quasicrystalline phases in the region of stronger hard-core repulsion of both hole-like mesophases. However, a full integration of these simulation results for quasicrystalline phases with the rest of the phase diagram is beyond the scope of the present study and will be undertaken in a future work.
The variety of mesophases found in this work highlights a general principle for systems with competing repulsive interactions at different length scales. While the specific locations of the phase boundaries depend on the details of the pair potential, the overall morphology of the phase diagram and the typical sequence of cluster crystals, stripes, and hole-like phases beneath the reentrant triangular crystal phase are expected to be qualitatively universal.
The detailed description of the short-wavelength region of the ground-state phase diagram presented can establish the basis for future investigations of thermal and quantum fluctuation effects on mesophases. For example, this system could be an excellent candidate for studying the melting of anisotropic two-dimensional crystalline phases [73, 74], including the possibility of intermediate phases between the ground-state and liquid phases [75], or, in the context of interacting bosons, for characterizing the corresponding quantum ground states [7, 32, 33, 76, 35]. Moreover, the classical behavior described here could be extended to higher densities, where longer-wavelength mesophases [3] may give rise to even more complex phenomena than those reported in the present study. Finally, we hope this study will stimulate further experimental investigations in colloidal systems under 2D confinement [21, 22].
Acknowledgements
We thank the computational resources provided by the Laboratório Multiusuário de Pesquisas Físicas (LAMPEF/FSC/CFM/UFSC) and by the Centro Nacional de Processamento de Alto Desempenho em São Paulo (CENAPAD-SP).